# A Survey on Human-AI Teaming with Large Pre-Trained Models

### v1

@@ Abstract
In the rapidly evolving landscape of artificial intelligence (AI), the collaboration between human intelligence and AI systems, known as Human-AI (HAI) Teaming, has emerged as a cornerstone for advancing problem-solving and decision-making processes. The advent of Large Pre-trained Models (LPtM) has significantly transformed this landscape, offering unprecedented capabilities by leveraging vast amounts of data to understand and predict complex patterns. This paper surveys the pivotal integration of LPtMs with HAI, emphasizing how these models enhance collaborative intelligence beyond traditional approaches. It examines the synergistic potential of LPtMs in augmenting human capabilities, discussing this collaboration for AI model improvements, effective teaming, ethical considerations, and their broad applied implications in various sectors. Through this exploration, the study sheds light on the transformative impact of LPtM-enhanced HAI Teaming, providing insights for future research, policy development, and strategic implementations aimed at harnessing the full potential of this collaboration for research and societal benefit.
1 Introduction
@@
The journey towards blending human intelligence with advanced technology spans several centuries, beginning with inventive attempts to simulate human intellect in machines. One notable example from the 1770s is the "Mechanical Turk," a machine that appeared to play chess autonomously but was, in fact, operated by a person concealed inside it [32]. This early endeavor, while not "Artificial Intelligence" (AI) in the modern sense, reflects the enduring fascination with creating technology that can mimic or complement human cognitive abilities. The formal realization of integrating human-like intelligence into machines culminated in 1956 at Dartmouth College, marking the official birth of Artificial Intelligence as a research field [134]. As AI technology has advanced and become more prevalent over the last decade, researchers have identified limitations within purely automated systems [110, 52, 133] leading to renewed focus on augmenting AI with human expertise, aiming to harness the best of both worlds (Fig. 0(a)). This approach seeks to enhance AI applications and foster a symbiotic relationship between human and artificial intelligence.
@@
Figure 1: A graphical depiction of the number of relevant articles submitted from 2013-2023. (a) illustrates the increasing engagement of research communities in collaborations between humans and AI over the years. (b) observes a notably sharp increase in submissions related to human-AI and large-scale models attributed to the recent emergence of large models.
@@
Navigating the rapidly advancing AI landscape, the integration of human cognition with Large Pre-trained Models (LPtM), including Large Language Models (LLM) [14, 151] and Large Vision Models (LVM) [80, 129, 69], has initiated a transformative shift. These models are "pre-trained" on vast datasets before being fine-tuned for specific tasks. This paradigm has opened new avenues for collaborative problem-solving and decision-making. LPtMs are shaped and refined by human expertise in terms of ethical guidance, creativity, and contextual understanding. In turn, AI amplifies human capabilities by processing data at large scales, offering insights and augmenting decision-making. This paper undertakes an extensive survey of Human-AI teaming, analyzing the complex interactions between human agents and sophisticated pre-trained models across various fields. Our study aims to uncover the progress made, confront the challenges, and understand the implications of this evolving partnership.
@@
1.1 Scope of the Survey
@@
In order to show support for the theme of our survey article, we try to implement human-AI teaming with the help of large language models to prepare this review. For the human part, the authors survey the existing literature about human-AI teaming, collect a subset of articles according to the desired structure, organize the survey into relevant sections, study submission statistics and put together tables and visualizations, and finally, prepare the first complete draft of the manuscript. Using the AI part in human-AI teaming, we ask ChatGPT-4 [120] for its feedback regarding the language flow of each subsection, placement of the citations, and revising the language wherever necessary. This helps in the ease in identifying the missing elements and making necessary revisions. The final manuscript is, therefore, a product of human and AI collaboration.
Figure 2: A representation of the scope of this survey. (a) We cover four broad categories - Model improvements, Effective HAI systems, Safe and Trustworthy AI, and their Applications - in the human-AI domain, along with the recent developments in them with the help of LPtM. The overlap between the four petals illustrates the fact that the four categories are not mutually exclusive; the articles may belong to two or more categories at the same time. (b) The heatmap shows a general trend of the number of articles co-existing within subcategories, including the articles discussing large models within each subject area. The overlap between the applications is not analyzed, given their specific and varied nature. AI: Artificial Intelligence, HAI: Human-AI, UI/UX: User Interface/User Experience.
1.2 Outline of the Survey
@@
In each segment of our study, we commence by delineating traditional methodologies employed in human-AI collaboration, subsequently delving into the contributions of extensive pre-trained models in the fulfillment of these tasks. Section 2 explores the involvement of incorporating human expertise into the AI model training cycle, fostering a cooperative relationship between humans and AI in active learning scenarios, enhancing learning through human feedback, and involving human experts in the thorough evaluation of machine learning models [65, 138, 181]. Section 3 delves into understanding the scope of effective Human-AI joint systems, optimizing user interfaces (UI) and system architectures emerges as a crucial area. It covers a range of topics under innovative UI designs and system structures aimed at boosting efficiency and user engagement in collaborative settings [88, 44].
@@
A paramount concern in the realm of Human-AI collaboration is the safety, security, and trustworthiness of AI systems. Section 4 of our comprehensive analysis encompasses multiple dimensions: mitigating algorithmic biases to uphold fairness, assessing the impact of Human-AI collaboration on workers’ autonomy, well-being, and job satisfaction, examining economic repercussions on employment and wages, addressing data privacy and security concerns, building trust in AI systems, and navigating the legal and regulatory frameworks governing Human-AI interactions [82, 62, 13]. These factors are integral to creating an ethical and responsible environment for AI’s integration into human-centric workflows. Furthermore, section 5 survey explores the broad spectrum of Human-AI teaming applications across various sectors. We investigate the unique challenges and opportunities in fields such as healthcare [11], autonomous vehicles, surveillance systems [60], gaming [46], education, and accessibility. Understanding the specific characteristics and benefits of Human-AI collaboration in these areas is key to influencing its future direction and maximizing its societal impact.
2 AI Model Improvements with Human-AI Teaming
@@
In this section, we delve into the process of training AI models, a critical stage that underpins their effectiveness and efficiency. Drawing upon the framework presented in Maadi et al. [100], we explore the integral role of human-AI interactions in model training. This framework delineates three interdependent phases: 1) Data control, where data is pre-processed and produced under stringent quality standards; 2) Machine Learning (ML) modeling and execution, guided by human expertise; and 3) ML evaluation and refinement. Our discussion primarily centers on the second phase, highlighting its pivotal role in shaping models that fulfill human-centric goals of effectiveness and efficiency. This section covers four key approaches: Human in the Loop, where human judgment directly influences AI training; Active Learning, focusing on AI systems that efficiently incorporate human input; Reinforcement Learning from Human Feedback, where AI behavior is shaped by human feedback; and Human Evaluated Machine Learning, emphasizing the role of human evaluation in refining AI models. Together, these subsections provide insights into how human-AI collaboration can enhance the training process, leading to more effective and reliable AI models.
@@
Human-AI Topics Subtopics Articles Cited
Section 2: AI model improvements with human-AI teaming
Human in the loop [169], [65], [105], [47], [138], [127], [95], [8], [90], [175], [29], [67], [87], [73], [144], [121], [181], [113], [130], [179], [72], [109], [177], [103]
Human evaluation in AI [6], [180], [5], [25], [65], [146], [34], [30], [121], [96], [24], [26], [152]
Section 3: Effective human-AI joint systems
Improving user interfaces for effective teaming [179], [60], [15], [128], [108], [172], [84], [135], [178], [25], [31], [160], [154], [45], [44], [164], [88]
Effective human-AI collaboration [164], [172], [126], [39], [93], [115], [178], [112], [53], [139], [131], [113], [121], [181], [180], [63], [103], [97], [162], [157], [123], [105], [65], [1], [48], [7], [21], [154], [94], [114], [116], [107], [75], [132], [140], [20]
Compatibility of human-AI systems [165], [25], [9], [162], [41], [58], [117], [103], [98], [64], [7], [171], [157], [63], [94], [112], [105], [22]
Section 4: Safe, secure and trustworthy AI
Algorithmic bias and fairness [54], [143], [33], [25], [111], [81], [66], [106], [83], [92], [50], [119], [10]
@@
Worker autonomy and well being [25], [82], [123], [37], [162], [170]
Effect on wages and jobs [62], [28], [25], [123], [3], [37], [162], [158]
Data privacy and security [43], [176], [74], [136], [89], [57], [153], [56], [36], [149], [137]
Trustworthy AI and accountability [64], [126], [18], [162], [79], [6]
Law and public policy [70], [19], [143], [12], [148], [159], [101]
Section 5: Applications
Healthcare [63], [11], [104], [102], [16], [27], [17], [99], [86], [71]
Autonomous vehicles [4], [97], [180], [48], [174], [124], [35], [167], [91]
Surveillance and security [125], [78], [60], [55], [23], [68]
Games [46], [173], [142], [2], [150], [156], [46], [65]
Education [118], [155], [168], [76], [40], [59], [38], [141], [42], [161], [61], [22]
Accessibility [85], [122], [77], [166], [51], [49], [147]
@@
Table 1: A tabular representation of the four broader focus topics covered in this survey, with their relevant subtopics. Each category contains its cited articles for the ease of reader reference.
2.1 Human in the Loop
@@
In the ‘Human in the Loop’ (HITL) approach, we explore the synergistic collaboration between human intellect and AI capabilities. This collaboration is especially critical in addressing AI’s limitations in self-learning and humans’ challenges in processing vast data within time constraints. The interplay of human and AI strengths leads to more robust and ethically sound AI systems. For instance, during the model design phase, human insights on ethical considerations, as highlighted by Whittlestone et al. [169], enrich AI perspectives beyond mere technical aspects. Conversely, in the model training and execution phase, the directional flow of information between humans and AI plays a pivotal role. Strategies such as InstructRL, as proposed by Hu and Sadigh [65], demonstrate the effectiveness of NLP instructions from humans in guiding AI, enhancing the model’s explainability and alignment with our expectations. Similarly, a research on AI-assisted brainstorming illustrates how AI’s engagement with human ideas can foster greater creativity and innovation [105].
@@
2.1.1 Human-AI Collaboration with Active Learning

### v2

@@ Abstract.
In the rapidly evolving landscape of artificial intelligence (AI), the collaboration between human intelligence and AI systems, known as Human-AI (HAI) Teaming, has emerged as a cornerstone for advancing problem-solving and decision-making processes. The advent of Large Pre-trained Models (LPtM) has significantly transformed this landscape, offering unprecedented capabilities by leveraging vast amounts of data to understand and predict complex patterns. This paper surveys the pivotal integration of LPtMs with HAI, emphasizing how these models enhance collaborative intelligence beyond traditional approaches. It examines the potential of LPtMs in augmenting human capabilities, discussing this collaboration for AI model improvements, effective teaming, ethical considerations, and their broad applied implications in various sectors. Through this exploration, the study sheds light on the transformative impact of LPtM-enhanced HAI Teaming, providing insights for future research, policy development, and strategic implementations aimed at harnessing the full potential of this collaboration for research and societal benefit.
Human-AI Teaming, Artificial Intelligence, Large Pre-Trained Models, Large Language Models
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Computing methodologies Artificial intelligenceccs: Human-centered computing Human computer interaction (HCI)
1. Introduction
@@
The journey towards blending human intelligence with advanced technology spans several centuries, beginning with inventive attempts to simulate human intellect in machines. One notable example from the 1770s is the "Mechanical Turk," a machine that appeared to play chess autonomously but was, in fact, operated by a person concealed inside it (Clark et al., 1999). This early endeavor, while not "Artificial Intelligence" (AI) in the modern sense, reflects the enduring fascination with creating technology that can mimic or complement human cognitive abilities. The formal realization of integrating human-like intelligence into machines culminated in 1956 at Dartmouth College, marking the official birth of Artificial Intelligence as a research field (Russell et al., 2021). As AI technology has advanced and become more prevalent over the last decade, researchers have identified limitations within purely automated systems (Mittelstadt et al., 2016; Goodfellow et al., 2015; Rodrigues, 2020) leading to renewed focus on augmenting AI with human expertise, aiming to harness the best of both worlds (Fig. 1(a)). This approach seeks to enhance AI applications and foster a symbiotic relationship between human and artificial intelligence.
@@
Figure 1. A graphical depiction of the number of relevant articles submitted from 2013-2023. (a) illustrates the increasing engagement of research communities in collaborations between humans and AI over the years. (b) observes a notably sharp increase in submissions related to human-AI and large-scale models attributed to the recent emergence of large models.
@@
Navigating the rapidly advancing AI landscape, the integration of human cognition with Large Pre-trained Models (LPtM), including Large Language Models (LLM) (Brown et al., 2020; Touvron et al., 2023) and Large Vision Models (LVM) (Kirillov et al., 2023; Radford et al., 2021; Jia et al., 2022), has initiated a transformative shift. These models are "pre-trained" on vast datasets before being fine-tuned for specific tasks. This paradigm has opened new avenues for collaborative problem-solving and decision-making. LPtMs are shaped and refined by human expertise in terms of ethical guidance, creativity, and contextual understanding. In turn, AI amplifies human capabilities by processing data at large scales, offering insights and augmenting decision-making. This paper undertakes an extensive survey of Human-AI teaming, analyzing the complex interactions between human agents and sophisticated pre-trained models across various fields. Our study aims to uncover the progress made, confront the challenges, and understand the implications of this evolving partnership.
@@
Figure 2. A representation of the scope of this survey. (a) We cover four broad categories - Model improvements, Effective HAI systems, Safe and Trustworthy AI, and their Applications - in the human-AI domain, along with the recent developments in them with the help of LPtM. The overlap between the four petals illustrates the fact that the four categories are not mutually exclusive; the articles may belong to two or more categories at the same time. (b) The heatmap shows a general trend of the number of articles co-existing within subcategories, including the articles discussing large models within each subject area. The overlap between the applications is not analyzed, given their specific and varied nature. AI: Artificial Intelligence, HAI: Human-AI, UI/UX: User Interface/User Experience.
1.1. Scope of the Survey
@@
In order to show support for the theme of our survey article, we try to implement human-AI teaming with the help of large language models to prepare this review. For the human part, the authors survey the existing literature about human-AI teaming, collect a subset of articles according to the desired structure, organize the survey into relevant sections, study submission statistics and put together tables and visualizations, and finally, prepare the first complete draft of the manuscript. Using the AI part in human-AI teaming, we ask ChatGPT-4 (OpenAI, 2023) for its feedback regarding the language flow of each subsection, placement of the citations, and revising the language wherever necessary. This helps in the ease in identifying the missing elements and making necessary revisions. The final manuscript is, therefore, a product of human and AI collaboration.
1.2. Outline of the Survey
@@
In each segment of our study, we commence by delineating traditional methodologies employed in human-AI collaboration, subsequently delving into the contributions of extensive pre-trained models in the fulfillment of these tasks. Section 2 explores the involvement of incorporating human expertise into the AI model training cycle, fostering a cooperative relationship between humans and AI in active learning scenarios, enhancing learning through human feedback, and involving human experts in the thorough evaluation of machine learning models (Hu and Sadigh, 2023; Settles, 2009; Ziegler et al., 2019). Section 3 delves into understanding the scope of effective Human-AI joint systems, optimizing user interfaces (UI) and system architectures emerges as a crucial area. It covers a range of topics under innovative UI designs and system structures aimed at boosting efficiency and user engagement in collaborative settings (Lemaignan et al., 2017; Fan et al., 2022).
@@
A paramount concern in the realm of Human-AI collaboration is the safety, security, and trustworthiness of AI systems. Section 4 of our comprehensive analysis encompasses multiple dimensions: mitigating algorithmic biases to uphold fairness, assessing the impact of Human-AI collaboration on workers’ autonomy, well-being, and job satisfaction, examining economic repercussions on employment and wages, addressing data privacy and security concerns, building trust in AI systems, and navigating the legal and regulatory frameworks governing Human-AI interactions (Konstantis et al., 2023; Hemmer et al., 2023; Braun et al., 2021). These factors are integral to creating an ethical and responsible environment for AI’s integration into human-centric workflows. Furthermore, section 5 survey explores the broad spectrum of Human-AI teaming applications across various sectors. We investigate the unique challenges and opportunities in fields such as healthcare (Bienefeld et al., 2023), autonomous vehicles, surveillance systems (Hauptman et al., 2023), gaming (Frans, 2021), education, and accessibility. Understanding the specific characteristics and benefits of Human-AI collaboration in these areas is key to influencing its future direction and maximizing its societal impact.
2. AI Model Improvements with Human-AI Teaming
@@
In this section, we delve into the process of training AI models, a critical stage that underpins their effectiveness and efficiency. Drawing upon the framework presented in Maadi et al. (Maadi et al., 2021), we explore the integral role of human-AI interactions in model training. This framework delineates three interdependent phases: 1) Data control, where data is pre-processed and produced under stringent quality standards; 2) Machine Learning (ML) modeling and execution, guided by human expertise; and 3) ML evaluation and refinement. Our discussion primarily centers on the second phase, highlighting its pivotal role in shaping models that fulfill human-centric goals of effectiveness and efficiency. This section covers four key approaches: Human in the Loop, where human judgment directly influences AI training; Active Learning, focusing on AI systems that efficiently incorporate human input; Reinforcement Learning from Human Feedback, where AI behavior is shaped by human feedback; and Human Evaluated Machine Learning, emphasizing the role of human evaluation in refining AI models. Together, these subsections provide insights into how human-AI collaboration can enhance the training process, leading to more effective and reliable AI models.
{tblr}
@@
width = colspec = Q[m,250]Q[m,350]Q[700], row1 = c, column2 = m, column3 = t, cell21 = r=2c, cell41 = r=3c, cell71 = r=6c, cell131 = r=6c, vlines, hline1,19 = -0.08em, hline2,4,7,13 = -, hline3,5-6,8-12,14-18 = 2-3, Human-AI Topics & Subtopics Articles Cited
Section 2: AI model improvements with human-AI teaming Human in the loop (Whittlestone et al., 2019), (Hu and Sadigh, 2023), (Memmert and Tavanapour, 2023), (Freeman et al., 2014), (Settles, 2009), (Pries et al., 2023), (Lu et al., 2022), (Bao et al., 2018), (Lertvittayakumjorn et al., 2020), (Yao et al., 2023), (Christiano et al., 2017), (Ibarz et al., 2018), (Lee et al., 2023), (Kaelbling et al., 1996), (Sutton et al., 1999), (Ouyang et al., 2022), (Ziegler et al., 2019), (Nakano et al., 2021), (Rafailov et al., 2023a), (Zhang et al., 2021), (Johnson and Vera, 2019), (Mirchandani et al., 2023), (Zha et al., 2023), (McNeese et al., 2021)
Human evaluation in AI (Bansal et al., 2019), (Zhou and Chen, 2019), (AV et al., 2022), (Chhibber et al., 2022), (Hu and Sadigh, 2023), (Tabrez et al., 2020), (Correia et al., 2023), (Clark et al., 2021), (Ouyang et al., 2022), (Lu et al., 2023), (Chen et al., 2024), (Chiang and Lee, 2023), (Uchendu et al., 2023)
Section 3: Effective human-AI joint systems Improving user interfaces for effective teaming (Zhang et al., 2021), (Hauptman et al., 2023), (Bryan Wang, 2023), (Qingxiao et al., 2023), (Mingming et al., 2023), (Wu et al., 2022), (Kozlowski and Ilgen, 2006), (Salas et al., 2017), (Zhang et al., 2023), (Chhibber et al., 2022), (Clark et al., 2016), (Wang et al., 2019), (van den Bosch et al., 2019), (Flathmann et al., 2021), (Fan et al., 2022), (Wei et al., 2023), (Lemaignan et al., 2017)
Effective human-AI collaboration (Wei et al., 2023), (Wu et al., 2022), (Pflanzer et al., 2022), (Dubey et al., 2020), (Liu et al., 2023a), (Ngo et al., 2023), (Zhang et al., 2023), (Munyaka et al., 2023), (Gopinath et al., 2022), (Sharifi Noorian et al., 2022), (Rafailov et al., 2023b), (Nakano et al., 2021), (Ouyang et al., 2022), (Ziegler et al., 2019), (Zhou and Chen, 2019), (Henry et al., 2022), (McNeese et al., 2021), (Lv et al., 2021), (Wang, 2023), (Veselovsky et al., 2023), (Pal, 2023), (Memmert and Tavanapour, 2023), (Hu and Sadigh, 2023), (Ahmad et al., 2023), (Fuchs et al., 2023), (Bansal et al., 2021), (Chakraborti et al., 2017), (van den Bosch et al., 2019), (Liu et al., 2023b), (Neerincx et al., 2018), (Nikolaidis et al., 2017), (Michelle Zhao, 2022), (Kaptein et al., 2016), (Rastogi et al., 2023), (Sharma et al., 2023), (Chakrabarty et al., 2023)
Compatibility of human-AI systems (Weisz et al., 2021), (Chhibber et al., 2022), (Bau et al., 2019), (Wang, 2023), (Eloundou et al., 2023), (Harrer, 2023), (Niu et al., 2020), (McNeese et al., 2021), (Lyons et al., 2019), (Hou et al., 2023), (Bansal et al., 2021), (Wu et al., 2022), (Veselovsky et al., 2023), (Henry et al., 2022), (Liu et al., 2023b), (Munyaka et al., 2023), (Memmert and Tavanapour, 2023), (Chang et al., 2023)
Section 4: Safe, secure and trustworthy AI Algorithmic bias and fairness (Green and Chen, 2019), (Stahl et al., 2021), (Conlon et al., 2023), (Chhibber et al., 2022), (Morrison et al., 2023), (Kirk et al., 2021), (Huang et al., 2023), (Meyer et al., 2023), (Kotek et al., 2023), (Li et al., 2023), (Gallegos et al., 2023), (Ohi et al., 2024), (Bi et al., 2023)
@@
Worker autonomy and well being (Chhibber et al., 2022), (Konstantis et al., 2023), (Pal, 2023), (Despotovic and Bogodistov, 2024), (Wang, 2023), (Woodruff et al., 2023)
Effect on wages and jobs (Hemmer et al., 2023), (Chowdhury et al., 2022), (Chhibber et al., 2022), (Pal, 2023), (Ashktorab et al., 2020), (Despotovic and Bogodistov, 2024), (Wang, 2023), (Walkowiak and MacDonald, 2023)
Data privacy and security (Ezer et al., 2019), (Yin et al., 2019), (Kaissis et al., 2020), (Samyuktha et al., 2022), (Lepri et al., 2021), (Hacker et al., 2023), (Ullah et al., 2023), (Gupta et al., 2023), (De Angelis et al., 2023), (Thapa and Adhikari, 2023), (Sebastian, 2023)
Trustworthy AI and accountability (Hou et al., 2023), (Pflanzer et al., 2022), (Caldwell et al., 2022), (Wang, 2023), (Kim et al., 2023), (Bansal et al., 2019)
Law and public policy (Jobin et al., 2019), (Cath et al., 2018), (Stahl et al., 2021), (Binz et al., 2023), (ThankGod Chinonso, 2023), (Walsh, 2023), (Marcos and Pullin, 2023)
Section 5: Applications Healthcare (Henry et al., 2022), (Bienefeld et al., 2023), (Memmert and Bittner, 2022), (McKinney et al., 2020), (Budd et al., 2021), (Choudhury and Asan, 2022), (Cai et al., 2019), (Lyu et al., 2023), (Kung et al., 2023), (Johnson et al., 2023)
Autonomous vehicles (Atakishiyev et al., 2021), (Lv et al., 2021), (Zhou and Chen, 2019), (Fuchs et al., 2023), (Yang et al., 2024), (Park et al., 2024), (Cui et al., 2024), (Wen et al., 2023), (Wang et al., 2023a), (Wang et al., 2023b)
Surveillance and security (Pazho et al., 2023), (Killcrece et al., 2003), (Hauptman et al., 2023), (Guo et al., 2018), (Chen et al., 2023), (Jain, 2023)
Games (Frans, 2021), (Xu et al., 2023), (Sobieszek and Price, 2022), (Akata et al., 2023), (Todd et al., 2023), (Värtinen et al., 2022), (Frans, 2021), (Hu and Sadigh, 2023)
Education (Nwana, 1990), (VanLehn, 2011), (Whitaker et al., 2013), (Kasinathan et al., 2017), (Eicher et al., 2018), (Hartle and Kaczorowski, 2019), (Dikli, 2006), (Silvia Milano, 2023), (Extance, 2023), (Wang and Demszky, 2023), (Hellas et al., 2023), (Chang et al., 2023)
Accessibility (Kumar and Jain, 2022), (Ozarkar et al., 2020), (Khan et al., 2020), (Wen et al., 2021), (Ghazal et al., 2021), (Gadiraju et al., 2023), (Taheri et al., 2023)
@@
Table 1. A tabular representation of the four broader focus topics covered in this survey, with their relevant subtopics. Each category contains its cited articles for the ease of reader reference.
2.1. Human in the Loop
@@
In the ‘Human in the Loop’ (HITL) approach, we explore the synergistic collaboration between human intellect and AI capabilities. This collaboration is especially critical in addressing AI’s limitations in self-learning and humans’ challenges in processing vast data within time constraints. The interplay of human and AI strengths leads to more robust and ethically sound AI systems. For instance, during the model design phase, human insights on ethical considerations, as highlighted by Whittlestone et al. (Whittlestone et al., 2019), enrich AI perspectives beyond mere technical aspects. Conversely, in the model training and execution phase, the directional flow of information between humans and AI plays a pivotal role. Strategies such as InstructRL, as proposed by Hu and Sadigh (Hu and Sadigh, 2023), demonstrate the effectiveness of NLP instructions from humans in guiding AI, enhancing the model’s explainability and alignment with our expectations. Similarly, a research on AI-assisted brainstorming illustrates how AI’s engagement with human ideas can foster greater creativity and innovation (Memmert and Tavanapour, 2023).
@@
2.1.1. Human-AI Collaboration with Active Learning